Roopa Patel won a prize for her work. She invested part of the money in a certificate of deposit at 4% and $3000 more than that amount in a bond paying 6%. Her annual interest income was %780. How much did Roopa invest at each rate?

Answer by checkley77(12844)

.04x+.06*(x+3,000)=780
.04x+.06x+180=780
.10x=780-180
.10x=600
x=600/.10
x=6,000 ans. for the amount invested @ 4%.
6,000+3,000=9,000 ans. for the amount invested @ 6%.
Proof:
.04*6,000+.06*9,000=780
240+540=780
780=780